# MultiVox: Benchmarking Voice Assistants for Multimodal Interactions

### v1

@@ 5.2 What causes these errors?
We conducted a manual error analysis of Gemini’s outputs on MultiVox to identify underlying failure patterns. Fig. 5 shows that perception errors dominate, accounting for 68.3% of all failures, primarily reflecting the model’s inability to ground to speech cues. Reasoning failures constitute 15.6% of errors, indicating that even when the model successfully perceives multimodal inputs, it struggles to effectively integrate this information to generate appropriate responses. Instruction understanding failures represent a similarly significant error category, where the model defaults to describing visual content rather than addressing the intended query. The overwhelming majority of speech perception errors indicates a clear bottleneck: improving speech perception capabilities is essential for building effective OVAs in multimodal contexts.
@@ 5.3 Do models really listen to speech cues?
To assess whether models are truly leveraging speech cues, we analyze their responses across confounder pairs, focusing on the top two performers: Gemini and Qwen2.5-Omni (Fig 5). Without considering confounders, both models appear moderately accurate—getting around 50% of responses correct (Gemini: 47%, Qwen: 51%) across these pairs. However, this aggregate accuracy can be misleading. When we examine whether models actually change their answers in response to flipped speech cues, we find that in over 45% of the confounder pairs, the model outputs are identical, i.e., they show NO FLIP in answer. Crucially, within those NO FLIP cases, most instances with one correct answer (77%) suggest that the correctness arises from chance or visual/textual bias—not from grounding in speech. This indicates that, despite non-trivial accuracy, models are largely ignoring non-verbal speech cues when answering.
@@ 6 Conclusion
We introduce MultiVox, the first benchmark designed to evaluate Omni Language Models (OLMs) as omni-modal voice assistants (OVAs) that integrate speech and vision for context-aware reasoning. Unlike prior benchmarks that rely on synthetic speech or focus only on unimodal cues, MultiVox includes 1000 professionally recorded, human-spoken questions paired with images or videos, emphasizing paralinguistic signals like tone, emotion, and background noise. A key innovation is the use of confounder pairs—speech variants with identical text and visuals—to ensure models attend to speech beyond surface cues. MultiVox enables fine-grained diagnosis across speech, vision, and general reasoning skills. Evaluation of 9 state-of-the-art models shows that, while visual grounding is robust, speech grounding remains a significant bottleneck. Our benchmark will be open-sourced to support the development of truly multimodal voice assistants.

### v2

@@ 5.2 What causes these errors?
We conducted a manual error analysis of Gemini’s outputs on MultiVox to identify underlying failure patterns. Fig. 5 shows that perception errors dominate, accounting for 75.3% of all failures, primarily reflecting the model’s inability to ground to speech cues. Reasoning failures constitute 7.0% of errors, indicating that even when the model successfully perceives multimodal inputs, it struggles to effectively integrate this information to generate appropriate responses. Instruction understanding failures represent a similarly significant error category, where the model defaults to describing visual content rather than addressing the intended query. The overwhelming majority of speech perception errors indicates a clear bottleneck: improving speech perception capabilities is essential for building effective OVAs in multimodal contexts.
@@ 5.3 Do models really listen to speech cues?
To assess whether models are truly leveraging speech cues, we analyze their responses across confounder pairs, focusing on the top two performers: Gemini 2.5 Pro and Qwen2.5-Omni (Fig 5). Without considering confounders, both models appear moderately accurate, getting around 50% of responses correct across these pairs. However, this aggregate accuracy can be misleading. When we examine whether models actually change their answers in response to flipped speech cues, we find that in a majority of the confounder pairs (57% for Gemini, 51% for Qwen), the model outputs are paraphrases, i.e., they show NO FLIP in answer. Crucially, within those NO FLIP cases, most instances with one correct answer suggest that the correctness arises from chance or visual/textual bias, not from grounding in speech. This indicates that, despite non-trivial accuracy, models are largely ignoring non-verbal speech cues when answering.
@@ 6 Conclusion
We introduce MultiVox, the first benchmark designed to evaluate Omni Language Models (OLMs) as omni-modal voice assistants (OVAs) that integrate speech and vision for context-aware reasoning. Unlike prior benchmarks that rely on synthetic speech or focus only on unimodal cues, MultiVox includes 1000 professionally recorded, human-spoken questions paired with images or videos, emphasizing paralinguistic signals like tone, emotion, and background noise. A key innovation is the use of confounder pairs—speech variants with identical text and visuals—to ensure models attend to speech beyond surface cues. MultiVox enables fine-grained diagnosis across speech, vision, and general reasoning skills. Evaluation of 10 state-of-the-art models shows that, while visual grounding is robust, speech grounding remains a significant bottleneck. Our benchmark will be open-sourced to support the development of truly multimodal voice assistants.
@@
Appendix A Appendix
In the Appendix, we provide:
  1. 1.
    Section B: Other Dataset Details
  2. 2.
    Section C: Annotator Details
  3. 3.
    Section D: LLM-as-a-judge Details
  4. 4.
    Section E: Voice Quality Assessment
Appendix B Other Dataset Details
Here we detail the categories and sub-categories present in our benchmark.
B.1 Acoustic Scene Understanding
Background Music Understanding: These tasks require the model to detect and interpret the presence and nature of background music in spoken queries. This includes genre classification, mood inference from music, and distinguishing music from speech or noise.
Sound Event Recognition: This task evaluates the model’s ability to detect and categorize discrete, identifiable audio events (e.g., dog barking, glass breaking, door closing) that occur within the auditory scene alongside spoken content.
Ambient Environment Sound: Models are tested on their capacity to recognize broader acoustic environments (e.g., airport, cafe, subway) based on background audio cues. We construct these scenes using the MS-Noise (Nachmani et al., 2024) dataset, overlaying clean speech with environmental recordings at a signal-to-noise ratio (SNR) of -2 dB to simulate challenging real-world conditions.
B.2 Paralanguage Understanding
Detailed distribution of the categories here are listed in Fig 7
Emotion: The task focuses on identifying the emotional state of the speaker as conveyed through prosodic features (e.g., pitch, energy), independent of lexical content.
Voice Modulation: Evaluates the model’s sensitivity to dynamic vocal variations such as emphasis, intonation, and expressiveness, which can affect intent or meaning.
Pronunciation: Tasks involve recognizing deviations from standard pronunciation, which may signal emotion, emphasis, or speaker background.
Volume: Assesses the ability to perceive and reason about loudness cues, which can convey urgency, emotion, or social context.
Pace: Evaluates how well the model understands speech rate—e.g., rushed versus slow delivery—as a cue to emotional or cognitive states.
Stuttering: Measures recognition and interpretation of speech disfluencies, including repeated sounds or syllables, as part of speaker modeling or intent understanding.
Breathiness: Focuses on detecting breathy voice quality, which may indicate fatigue, emotion, or affective state.
B.3 Speaker Profiling
Biological Gender Estimation: The task is to estimate the speaker’s biological gender based solely on voice characteristics, controlling for content and visual input.
Age Group Estimation: Models must infer the age group (e.g., child, adult, elderly) of the speaker using acoustic cues.
Appendix C Annotator Details
1. Annotator Composition
We formed a panel of six domain experts for our dataset creation and filtering process and our dataset review process. The panel consisted of four Ph.D students pursuing speech and audio-visual research and two MS students having research in speech and audio processing. The expertise of all domain experts is evidenced by their research publications and contributions to the field.
2. Meetings to decide question creation process
All six annotators had three 2-hour online meetings to discuss the question and corresponding answers creation process to reach a consensus of the pipeline to be followed for dataset creation. The online meetings covered these details:
  • Multimodal Question-Answering Foundations: Aligning nonverbal audio cues (e.g., tone, background sounds) with visual context (e.g., scene imagery).
  • Confounder Pair Design: Generating minimal audio variants that invert answers (e.g., adding subtle noise to flip “quiet” vs. “noisy”).
  • Annotation Platform & Guidelines: Hands-on use of our custom interface and details about input/output formats, edge cases, and scoring rubrics.
Following the online meetings, annotators jointly labeled 50 pilot samples and attained $\geq 90\%$ inter-annotator agreement before proceeding to full‐scale work.
3. Question Creation Process
Each annotator followed a three-step pipeline:
  • Scenario Drafting: Write a conversational prompt targeting modality cues (e.g., “Can you tell if this room is too loud for a conference call?”).
  • Confounder Generation: Create a paired version of the prompt with the same text and visuals but alter one audio attribute (e.g., add background machinery noise).
  • Media Pairing: Select or synthesize matching audio and visual assets from our curated libraries to illustrate both the original and confounded conditions.
This ensured every question/confounder pair isolated the intended cue and prevented shortcut learning by downstream models.
Figure 6: Pie charts showing different Speaker Profile categories in terms of Gender and Age
Figure 7: Pie charts showing different Paralanguage Understanding Categories and different Emotion Categories in the benchmark
4. Answer Creation Process
For each question instance, annotators produced:
  • Reference Answer: A concise response directly addressing the prompt (e.g., “No, it’s quiet enough for clear conversation”).
  • Rationale Statement: A brief explanation linking the critical cue to the answer (e.g., “The low ambient noise level confirms a silent office setting”).
These reference answers and rationales formed the ground truth for our GPT-4–based judgment of model responses.
5. Annotation Criteria & Quality Control
All questions and answers were crafted according to these overarching principles:
  • Clarity & Brevity: Simple, conversational language devoid of syntactic complexity.
  • Modality Isolation: Exactly one audio or visual “hook” per item, ensuring focused evaluation.
  • Balanced Distribution of Skills: Even distribution of questions across different skills.
Appendix D LLM-as-a-Judge Details
  1. 1.
    Human vs. LLM Evaluation Experiment: To validate whether a large language model (LLM) could reliably substitute for our expert reviewers, we conducted a blind evaluation on a sample of 300 question-answer pairs drawn evenly from our benchmark. Each pair was independently graded on a 1-5 scale by:
    • Three domain experts from our panel following a predecided grading criteria
    • An LLM-as-a-Judge, implemented via a single GPT-4.1-mini call per sample.
    We computed average scores for each item under both conditions and measured inter-rater agreement using Cohen’s $\kappa$. Across all 300 items, human–human agreement averaged $\kappa=0.82$, while human–LLM agreement reached $\kappa=0.78$, indicating the LLM’s judgments were almost as consistent with the experts’ as the experts were with one another. Overall the LLM’s ratings fell within one point of the experts’ in 92% of cases.
    Given these results, we adopted the LLM-as-a-Judge for large-scale scoring in subsequent experiments, leveraging its efficiency without materially sacrificing quality.
  2. 2.
    LLM-as-a-Judge Criteria:
    We present each question to the LLM-as-a-Judge by supplying both the corresponding speech and visual inputs. The model is then asked to rate the provided answer on a five-point scale according to the following criteria:
    • Score 1: The response is often off-topic or incorrect and fails to recognize or use the specified speaker characteristic.
    • Score 2: The response occasionally addresses the prompt but handles the speaker characteristic inconsistently or superficially.
    • Score 3: The response shows a basic understanding of intent, with partial integration of the speaker characteristic but lacks depth or precision.
    • Score 4: The response delivers relevant and mostly accurate content that usually incorporates the speaker characteristic, with only minor lapses.
    • Score 5: The response consistently produces accurate, context-rich answers that fully and effectively integrate the speaker characteristic.
Appendix E Voice Quality Assessment
We generated synthetic audio for 100 benchmark questions using two text-to-speech systems: CosyVoice (open source) (Du et al., 2024) and ElevenLabs (ElevenLabs Inc., 2025) (commercial). We also recorded these questions by professional voice actors.
To compare these renderings, we conducted a Mean Opinion Score (MOS) Test on a random subset of 100 paralanguage-focused queries (as in Sec. 3.4). Ten expert annotators rated each audio on two dimensions, using a 5-point scale:
  • Attribute match: How accurately the intended paralinguistic cue (e.g., emotion, background noise) was conveyed.
  • Naturalness: The overall human-likeness of the voice.
The MOS results were as follows: Professional human recordings: 4.6 (attribute match) / 4.5 (naturalness), CosyVoice TTS: 2.4 / 2.1, ElevenLabs TTS: 3.1 / 3.3 .
These findings confirm that, while modern TTS can approximate certain prosodic features, professional voice actors remain far superior in both fidelity and naturalness, which led us to use human-recorded queries by preofessional voice actors throughout the benchmark.
Appendix F Voice Data Collection
To ensure natural and expressive speech, we employed four professional voice actors, four male and two female, contracted via Fiverr. Each actor was compensated according to the standard rates listed on the platform. This approach allowed us to capture high-quality, emotionally varied recordings with realistic prosody and delivery across all tasks. Our institution’s Institutional Review Board (IRB) has granted approval for this data collection.
Appendix G Additional Details: Auxiliary
Compute Infrastructure: All our experiments are conducted on a single NVIDIA A6000 GPU. No training is required, and depending on the downstream task, a single inference run on a benchmark requires anywhere between 30 minutes to 2 hours.
Implementation Software and Packages: We use the official implementation of the OLMs we benchmark. For LLM-as-judge implementation, we utilize the official OpenAI APIs.
Potential Risks: We manually curate all our questions to avoid any potential harmful or biased samples.